OpenStudy (anonymous):

Hello I need halp with this 10th grade geometry question find the equation of the circle with the following center c(-2.1) with point P (4,9)on the circle

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Well, the radius of the circle will be the distance between those two points

OpenStudy (anonymous):

d^2 = (4 - (-2))^2 + (9-1)^2 = 6^2 + 8^2 = 100, so d= the radius = 10

OpenStudy (anonymous):

so the equation of the circle would be (x+2)^2 + (y-1)^2 = 100

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes it does thank you so very much

OpenStudy (anonymous):

no problem, any time!